OpenStudy (anonymous):

shawna can paint a fence in 5 hrs. Jassie can paint the same fence in 10 hrs. how long will it take them to paint the fence together56 seconds ago Live preview of your reply:shawna can paint a fence in 5 hrs. Jassie can paint the same fence in 10 hrs. how long will it take them to paint the fence together2 minutes ago

OpenStudy (watchmath):

shawna speed = 1/5 ( 1 unit of fence for every 5 hours) jessie speed = 1/10 combine speed = 1/5 + 1/10 = 3/10 So they can finish 3 fence for every 10 hours For one fence they just need 10/3 hours.